# Mixed Newton Method for Optimization in Complex Spaces

### v1

@@ 1 Introduction
$\frac{\partial^{2}f}{\partial\bar{z}\partial z}=\sum_{j}\overline{\frac{dg_{j}}{dz}}\frac{dg_{j}}{dz}^{\top}$
@@ 2 Mixed Newton method with regularization
Recall that we minimize a function $f(z)=\sum_{j}|g_{j}(z)|^{2}$ with all $g_{j}$ holomorphic functions. Set $B=\frac{\partial^{2}f}{\partial\bar{z}\partial z}=\sum_{j}\overline{g_{j}^{\prime}}(g_{j}^{\prime})^{\top}$, $A=\sum_{j}g_{j}\overline{g_{j}^{\prime\prime}}$, where the derivatives are evaluated at the critical point $\hat{z}$ satisfying $\frac{\partial f}{\partial\bar{z}}=\sum_{j}g_{j}\overline{g_{j}^{\prime}}=0_{n}$. Note that $B=B^{*}\succeq 0_{n\times n}$, $A=A^{\top}$, i.e., $B$ is complex-hermitian positive semi-definite, and $A$ is symmetric, but in general not hermitian.
@@
$\frac{\partial f(z)}{\partial\bar{z}}$ $=\sum_{j}g_{j}(z)\overline{g_{j}^{\prime}(z)}=\sum_{j}(g_{j}(\hat{z})+(g_{j}^{\prime}(\hat{z}))^{\top}\delta+O(\|\delta\|^{2}))(\overline{g_{j}^{\prime}(\hat{z})+g_{j}^{\prime\prime}(\hat{z})\delta+O(\|\delta\|^{2})})$
@@
$\frac{\partial^{2}f(z)}{\partial\bar{z}\partial z}$ $=\sum_{j}\overline{g_{j}^{\prime}(z)}(g_{j}^{\prime}(z))^{\top}=B+O(\|\delta\|).$
@@
Let now $H_{\mathbb{R}}=\frac{\partial^{2}f(\hat{z})}{\partial(\operatorname{Re}(z),\operatorname{Im}(z))^{2}}$ be the Hessian at the critical point in real coordinates $(\operatorname{Re}(z),\operatorname{Im}(z))$. Then the signature of $H_{\mathbb{R}}$ is responsible for the type of the critical point. If $H_{\mathbb{R}}\succ 0_{n\times n}$, then $\hat{z}$ is a minimum, and if the signature is mixed (with positive and negative eigenvalues), then the point $\hat{z}$ is a saddle point of the cost function $f$.
@@
The dynamical system defined by $Y^{\prime}$ is strictly stable if and only if its spectrum is contained in the interval $(-1,1)$. This is equivalent to the condition that the spectrum of $M^{\prime}=I_{n}-Y^{\prime}$ is contained in the interval $(0,2)$, in which case $M^{\prime}\succ 0_{n\times n}$. Hence stability of the dynamical system defined by the RMNM in the vicinity of a critical point implies that the point is a minimum. This still leaves open the possibility that some minima are repelling, namely, if the spectrum of $M^{\prime}$ is positive and its maximum eigenvalue exceeds 2.
Let us prove that this situation is not possible. Assume that $M^{\prime}\succ 0_{n\times n}$. Then the diagonal blocks of $M^{\prime}$ are also positive definite. In particular, we get $0_{n\times n}\prec Q\prec I_{n}$, where the first inequality holds by the definition of $Q$. Hence $Q$ and $I_{n}-Q$ are contracting, and $(I_{n}-Q)^{-1/2},(I_{n}-\bar{Q})^{-1/2}$ are expanding, i.e., application of these maps strictly increases the 2-norm of any non-zero vector. From $M^{\prime}\succ 0_{n\times n}$ it follows that $\sigma_{\max}((I_{n}-Q)^{-1/2}S(I_{n}-\bar{Q})^{-1/2})<1$, i.e., this matrix product is contracting. But then $S$ must be contracting, and all 4 blocks of the matrix $M^{\prime}$ are contracting. It follows that $\lambda_{\max}(M^{\prime})=\sigma_{\max}(M^{\prime})<2$. The case of repelling minima is hence excluded.
@@ 3 Minimization of real analytic functions
In this section, we study how the MNM can be used to minimize a real analytic function $F$ on a simply connected domain $D_{\mathbb{R}}\subset\mathbb{R}^{n}$. The idea is to extend the function to a domain $D_{\mathbb{C}}\subset\mathbb{C}^{n}$ in complex space, to yield a function of the form $|g|^{2}$, where $g:D_{\mathbb{C}}\to\mathbb{C}$ is a holomorphic function.
Clearly, this is only possible if $F>0$ on the domain $D_{\mathbb{R}}$. Since we may add constants to the objective, we may assume this condition without loss of generality. Then $F^{2}$ is also analytic on $D_{\mathbb{R}}$, moreover, minimizing $F^{2}$ is equivalent to minimizing $F$. If we define $g$ to be the holomorphic continuation of $F$ to a simply connected domain $D_{\mathbb{C}}$, then $F^{2}$ will be the restriction of a function $f:D_{\mathbb{C}}\to\mathbb{R}$ of the desired form $f=|g|^{2}$ with $g$ holomorphic.
This straightforward approach meets the following obstacles. The sum $\frac{\partial^{2}f}{\partial\bar{z}\partial z}=\sum_{j}\overline{\frac{dg_{j}}{dz}}\frac{dg_{j}}{dz}^{\top}$ defining the mixed Hessian consists of only one summand, hence the mixed Hessian is rank 1. Another problem is that the minima of $f$ can be located in the complex plane and a local minimum of $f$ on $D_{\mathbb{R}}$ may be a saddle point of $f$ on $D_{\mathbb{C}}$. The iterates will then be pushed away from $D_{\mathbb{R}}$.
@@
Finally, we obtain the following iteration of the MNM for minimization of a positive analytic function $F=g_{0}$ on $D_{\mathbb{R}}$ with additive complex-repulsive regularization:
@@ 4 Numerical experiments
In this section, we present the results of our experiments for the comparison of the RMNM (3) and the ONM (4). We consider the problem of minimization of positive real polynomials on $\mathbb{R}^{2}$.
@@
Functions (6), (7) and (8) are positive non-convex. For these functions, we run RMNM and ONM, with 625 different initial points in the square $[-1,2]\times[-1,2]$ for Example 1, with 1024 initial points in the square $[-1,3]\times[-1,3]$ for Example 2, and with different 2601 initial points in the square $[-3,2]\times[-3,2]$ for Example 3. The initial points for both methods are the same and real. The results are presented in Table 1 and Table 2, together with the used parameter value $\gamma$. These results show the number of initial points for which the methods converge to the local or global minimum and to the saddle point, they also show the number of initial points for which the methods diverge.
@@ 5 Conclusions
We compared the method on two training tasks for neural networks, one with application in telecommunications, the other a benchmark network. The regularized versions of MNM showed superior performance and required computational resources reduction in comparison with the classical competitors.
@@ A.1 Applications
$\displaystyle\text{MSE}_{\text{min}}=\min_{j}(\text{MSE}_{j}),\quad\displaystyle\text{MSE}_{\text{max}}=\min_{j}(\text{MSE}_{j}),\quad j=\overline{0,N-1},$ (9)
@@
$\displaystyle\text{MSE}_{\text{min}}=\min_{j,p}(\text{MSE}_{j,p}),\quad\displaystyle\text{MSE}_{\text{max}}=\min_{j,p}(\text{MSE}_{j,p}),\quad j=\overline{0,N-1},\quad p=\overline{0,P-1},$ (10)
@@ A.2 Derivation of the iterations of LM-MNM and CMNM
Define $\Delta_{c}=c_{1}-c\in\mathbb{C}^{2n}$ and $\Delta_{z}=z_{1}-z$, then $\|\Delta_{r}\|^{2}=\frac{1}{2}\|\Delta_{c}\|^{2}$. The Taylor polynomial can be rewritten as
@@ A.2.1 Digital Pre-Distortion
where $\mathbb{K}=\mathbb{R}$ in case of RV-CNN, $\mathbb{K}=\mathbb{C}$ in case of CV-CNN, $\sigma(\cdot)$ – sigmoid function. In addition, $c$ is the input channel index, $p$ the layer index, $n$ the time index, $z_{c,k,p}\in\mathbb{K}^{1\times M}$ the 1D convolution kernel of the input channel $c$, the output channel $k$, the layer $p$. Further $M$ is the kernel size, $b_{c,p}\in\mathbb{K}$ the bias, $u_{n,p-1,c}\in\mathbb{K}^{1\times L}$ the $c$-th output of convolutional layer $p-1$. Note that the input of the 00-th convolutional layer is formally given by the model input channel, $u_{n,-1,c}=x_{n,c}\in\mathbb{K}^{1\times L}$.

### v2

@@ 1 Introduction
$\frac{\partial^{2}f}{\partial\bar{z}\partial z}=\sum_{j=0}^{m-1}\overline{\frac{dg_{j}}{dz}}\frac{dg_{j}}{dz}^{\top}$
@@ 2 Mixed Newton method with regularization
Recall that we minimize a function $f(z)=\sum_{j=0}^{m-1}|g_{j}(z)|^{2}$ with all $g_{j}$ holomorphic functions. Set $B=\frac{\partial^{2}f}{\partial\bar{z}\partial z}=\sum_{j=0}^{m-1}\overline{g_{j}^{\prime}}(g_{j}^{\prime})^{\top}$, $A=\sum_{j=0}^{m-1}g_{j}\overline{g_{j}^{\prime\prime}}$, where the derivatives are evaluated at the critical point $\hat{z}$ satisfying $\frac{\partial f}{\partial\bar{z}}=\sum_{j=0}^{m-1}g_{j}\overline{g_{j}^{\prime}}=0_{n}$. Note that $B=B^{*}\succeq 0_{n\times n}$, $A=A^{\top}$, i.e., $B$ is complex-hermitian positive semi-definite, and $A$ is symmetric, but in general not hermitian.
@@
$\frac{\partial f(z)}{\partial\bar{z}}$ $=\sum_{j=0}^{m-1}g_{j}(z)\overline{g_{j}^{\prime}(z)}=\sum_{j=0}^{m-1}(g_{j}(\hat{z})+(g_{j}^{\prime}(\hat{z}))^{\top}\delta+O(\|\delta\|^{2}))(\overline{g_{j}^{\prime}(\hat{z})+g_{j}^{\prime\prime}(\hat{z})\delta+O(\|\delta\|^{2})})$
@@
$\frac{\partial^{2}f(z)}{\partial\bar{z}\partial z}$ $=\sum_{j=0}^{m-1}\overline{g_{j}^{\prime}(z)}(g_{j}^{\prime}(z))^{\top}=B+O(\|\delta\|).$
@@
Let now $H_{\mathbb{R}}=\frac{\partial^{2}f(\hat{z})}{\partial(\operatorname{Re}(z),\operatorname{Im}(z))^{2}}$ be the Hessian at the critical point in real coordinates $(\operatorname{Re}(z),\operatorname{Im}(z))$. Then the signature of $H_{\mathbb{R}}$ is responsible for the type of the critical point. If $H_{\mathbb{R}}\succ 0_{2n\times 2n}$, then $\hat{z}$ is a minimum, and if the signature is mixed (with positive and negative eigenvalues), then the point $\hat{z}$ is a saddle point of the cost function $f$.
@@
The dynamical system defined by $Y^{\prime}$ is strictly stable if and only if its spectrum is contained in the interval $(-1,1)$. This is equivalent to the condition that the spectrum of $M^{\prime}=I_{2n}-Y^{\prime}$ is contained in the interval $(0,2)$, in which case $M^{\prime}\succ 0_{2n\times 2n}$. Hence stability of the dynamical system defined by the RMNM in the vicinity of a critical point implies that the point is a minimum. This still leaves open the possibility that some minima are repelling, namely, if the spectrum of $M^{\prime}$ is positive and its maximum eigenvalue exceeds 2.
Let us prove that this situation is not possible. Assume that $M^{\prime}\succ 0_{2n\times 2n}$. Then the diagonal blocks of $M^{\prime}$ are also positive definite. In particular, we get $0_{n\times n}\prec Q\prec I_{n}$, where the first inequality holds by the definition of $Q$. Hence $Q$ and $I_{n}-Q$ are contracting, and $(I_{n}-Q)^{-1/2},(I_{n}-\bar{Q})^{-1/2}$ are expanding, i.e., application of these maps strictly increases the 2-norm of any non-zero vector. From $M^{\prime}\succ 0_{2n\times 2n}$ it follows that $\sigma_{\max}((I_{n}-Q)^{-1/2}S(I_{n}-\bar{Q})^{-1/2})<1$, i.e., this matrix product is contracting. But then $S$ must be contracting, and all 4 blocks of the matrix $M^{\prime}$ are contracting. It follows that $\lambda_{\max}(M^{\prime})=\sigma_{\max}(M^{\prime})<2$. The case of repelling minima is hence excluded.
@@ 3 Minimization of real analytic functions
In this section, we study how the MNM can be used to minimize a real analytic function $F$ on a simply connected domain $D_{\mathbb{R}}\subseteq\mathbb{R}^{n}$. The idea is to extend the function to a domain $D_{\mathbb{C}}\subseteq\mathbb{C}^{n}$ in complex space, to yield a function of the form $|g|^{2}$, where $g:D_{\mathbb{C}}\to\mathbb{C}$ is a holomorphic function.
Clearly, this is only possible if $F\geq 0$ on the domain $D_{\mathbb{R}}$. Since we may add constants to the objective, we may assume this condition without loss of generality. Then $F^{2}$ is also analytic on $D_{\mathbb{R}}$, moreover, minimizing $F^{2}$ is equivalent to minimizing $F$. If we define $g$ to be the holomorphic continuation of $F$ to a simply connected domain $D_{\mathbb{C}}$, then $F^{2}$ will be the restriction of a function $f:D_{\mathbb{C}}\to\mathbb{R}$ of the desired form $f=|g|^{2}$ with $g$ holomorphic.
This straightforward approach meets the following obstacles. The sum $\frac{\partial^{2}f}{\partial\bar{z}\partial z}=\sum_{j=0}^{m-1}\overline{\frac{dg_{j}}{dz}}\frac{dg_{j}}{dz}^{\top}$ defining the mixed Hessian consists of only one summand, hence the mixed Hessian is rank 1. Another problem is that the minima of $f$ can be located in the complex plane and a local minimum of $f$ on $D_{\mathbb{R}}$ may be a saddle point of $f$ on $D_{\mathbb{C}}$. The iterates will then be pushed away from $D_{\mathbb{R}}$.
@@
Finally, we obtain the following iteration of the MNM for minimization of a nonnegative analytic function $F=g_{0}$ on $D_{\mathbb{R}}$ with additive complex-repulsive regularization:
@@ 4 Numerical experiments
In this section, we present the results of our experiments for the comparison of the RMNM (3) and the ONM (4). We consider the problem of minimization of nonnegative real polynomials on $\mathbb{R}^{2}$.
@@
Functions (6), (7) and (8) are nonnegative and non-convex. For these functions, we run RMNM and ONM, with 625 different initial points in the square $[-1,2]\times[-1,2]$ for Example 1, with 1024 initial points in the square $[-1,3]\times[-1,3]$ for Example 2, and with different 2601 initial points in the square $[-3,2]\times[-3,2]$ for Example 3. The initial points for both methods are the same and real. The results are presented in Table 1 and Table 2, together with the used parameter value $\gamma$. These results show the number of initial points for which the methods converge to the local or global minimum and to the saddle point, they also show the number of initial points for which the methods diverge.
@@ 5 Conclusions
We compared the method on two training tasks for neural networks, one with application in telecommunications, the other a benchmark network. The regularized versions of MNM showed superior performance and required computational resources reduction in comparison with the classical competitors. These results are presented in the Appendix.
@@ A.1 Applications
$\displaystyle\text{MSE}_{\text{min}}=\min_{j}(\text{MSE}_{j}),\quad\displaystyle\text{MSE}_{\text{max}}=\max_{j}(\text{MSE}_{j}),\quad j=\overline{0,N-1},$ (9)
@@
$\displaystyle\text{MSE}_{\text{min}}=\min_{j,p}(\text{MSE}_{j,p}),\quad\displaystyle\text{MSE}_{\text{max}}=\max_{j,p}(\text{MSE}_{j,p}),\quad j=\overline{0,N-1},\quad p=\overline{0,P-1},$ (10)
@@ A.2 Derivation of the iterations of LM-MNM and CMNM
Define $\Delta_{c}=c_{1}-c\in\mathbb{C}^{2n}$ and $\Delta_{z}=z_{1}-z\in\mathbb{C}^{n}$, then $\|\Delta_{r}\|^{2}=\frac{1}{2}\|\Delta_{c}\|^{2}$. The Taylor polynomial can be rewritten as
@@ A.2.1 Digital Pre-Distortion
where $\mathbb{K}=\mathbb{R}$ in case of RV-CNN, $\mathbb{K}=\mathbb{C}$ in case of CV-CNN, $\sigma(\cdot)$ – sigmoid function. In addition, $c$ is the input channel index, $p$ the layer index, $n$ the time index, $z_{c,k,p}\in\mathbb{K}^{1\times M}$ the 1D convolution kernel of the input channel $c$, the output channel $k$, the layer $p$. Further $M$ is the kernel size, $b_{c,p}\in\mathbb{K}$ the bias, $u_{n,p-1,c}\in\mathbb{K}^{1\times L}$ the $c$-th output of convolutional layer $p-1$. Note that the input of the 00-th convolutional layer is formally given by the model input channel, $u_{n,-1,c}=\widetilde{x}_{n,c}\in\mathbb{K}^{1\times L}$.